# Towards Imperceptible Backdoor Attack in Self-supervised Learning

### v1

@@ C.1.2 Grad-CAM
We assess IMPERATIVE’s resilience to Grad-CAM [64], which generates a heatmap for a given model and input sample, where the heat intensity of each pixel reflects its significance in the model’s final prediction. Grad-CAM is effective for identifying smaller Trojans [45, 18], as these Trojans tend to produce high heat values in compact trigger zones, leading to abnormal heatmaps. Nevertheless, our backdoor transformation function alters the entire image, rendering Grad-CAM ineffective in detecting it. Figure 5 illustrates the comparison of heatmaps from both clean images and backdoored images created by our method. The similarity in these heatmaps suggests that IMPERATIVE can withstand defenses based on Grad-CAM.
C.1.3 Various Noises
@@
Next, we are going to introduce the noises, i.e., JPEG compression, Salt&Pepper noise, and Poisson noise, applied in the robustness test.
@@
Figure 6: Resilience to JPEG compression, Salt&Pepper noise, Poisson noise of different densities.
@@
Evaluating the robustness of our method across these noises helps ensure its performance is reliable even when image quality is compromised when encountering real-world image issues. We illustrate the evaluation results in Figure 6, where the pre-trained dataset is CIFAR10 and the downstream dataset is STL10. The results prove that our method, which utilizes $\mathcal{L}_{disentangle}$, maintains high ASRs in most cases, and the ASR is significantly higher than the cases ablating $\mathcal{L}_{disentangle}$ in our method. For example, under different intensities of Poisson noise, the ASR of our method maintains nearly 100%, while the ASR of the cases ablating $\mathcal{L}_{disentangle}$ in our method is unstable, ranging from 12% to 61%. The outcome indicates that the $\mathcal{L}_{disentangle}$ plays a prominent role in promoting the robustness of our method. The reason may be that the $\mathcal{L}_{disentangle}$ encourages the model to learn features that remain invariant under the transformations. Namely, the model can still recognize the embedded triggers even when various noises or other image perturbations are applied. This aspect of feature preservation is key to ensuring that the backdoor attack remains effective regardless of the image transformations or certain noises applied.
C.1.4 STRIP
Figure 7: Resilient to STRIP [16]. IMPERATIVE can withstand the STRIP runtime defense.
@@ C.1.4 STRIP
STRIP [16] scrutinizes a suspect sample by overlaying diverse image patterns onto it and monitoring the uniformity of the model’s predictions for these altered inputs. A low entropy score, suggesting consistent predictions across these perturbed samples, would lead STRIP to classify the sample as Trojan-infected. The experimental outcomes depicted in Figure 7 reveal that the entropy values for both clean and Trojan-compromised models produced through our method overlap considerably, signifying that our attack can withstand the STRIP runtime defense. IMPERATIVE manages to evade STRIP as the superimposition process destroys the attack’s trigger, thereby causing both the Trojan-induced and clean samples to exhibit significant alterations in prediction when superimposed, aligning them with what is observed in clean cases.
C.2 Sensitivity Evaluation
Figure 8: Sensitivity evaluation of hyperparameters. IMPERATIVE performs consistently over these parameter ranges.
@@ C.2 Sensitivity Evaluation
Figure 8 illustrates the sensitivity evaluation of the hyperparameters alpha and beta, showcasing that the model’s performance exhibits relative insensitivity to parameter variations. Specifically, for alpha, both the Attack Success Rate (ASR) and Backdoored Accuracy (BA) remain relatively constant across a range of alpha values from 0.05 to 0.30, suggesting that once an optimal threshold is reached, the model’s performance is stable despite further adjustments to alpha. Similarly, for beta, changes across a broader spectrum, from 0 to 10, show that ASR quickly stabilizes at 100%, while BA and CA display minimal fluctuations, indicating the robust model performance that is not significantly affected by variations in beta. The horizontal axis in both graphs represents the incremental values of alpha and beta, demonstrating the model’s consistent performance over these parameter ranges.
C.3 Results Analysis
@@
Figure 9: Visualization of feature space segregation over epochs demonstrates the efficacy of our backdoor trigger in creating a distinguishable separation between backdoor and clean features, unlike WaNet [51] where the features remain intermingled, highlighting our method’s utility in securing encoders against backdoor threats.
@@
Figure 9 presents a visual analysis of the feature shifts induced by backdoor injection across different epochs for our proposed method compared to the WaNet [51]. Throughout the backdoor injection process, it becomes evident that our method’s trigger effectively separates the backdoor features from the clean features. This distinction is observable as the training progresses from epoch 10 through to epoch 200, where the separation between the backdoor and clean features becomes increasingly pronounced, indicating a clear demarcation that our encoder can exploit to differentiate between the two.
@@
Appendix D Implementation Details
The following sections provide a comprehensive breakdown of the process of implementing the attack, which is methodically organized into three distinct phases: pre-training the encoder (Appendix D.1), pre-training the backdoor injector (Appendix D.2), injecting the backdoor (Appendix D.3), and training the downstream classifiers (Appendix D.4). Additionally, we explain the HSV&HSL Color Spaces and Wasserstein Distance (Appendix D.5) used in the method in detail. Moreover, we provide more details on the datasets (Appendix D.6.1), evaluation metrics (Appendix D.6.2), and experimental setup details (Appendix D.6.3, Appendix D.6.4).
D.1 Pre-training Image Encoders
@@
D.2 Pre-training Backdoor Injector
To enhance the optimization speed and increase the effectiveness of the backdoor attack, we use the method outlined in Algorithm 2 to pre-train a backdoor injector, which serves as the initial setup for the injector in Algorithm 1. We use the U-Net architecture [59] for the backdoor injector. The detailed architecture is shown in the Table 11.
Algorithm 2 Pre-training backdoor injector
1:function Pre-training injector($\mathcal{D}_{s}$)
3:     for $\text{iter}=0$ to max_epochs do
4:         ${\mathcal{D}_{s}}^{{}^{\prime}}\leftarrow\mathcal{I}_{\phi}(\mathcal{D}_{s})$
6:         $\mathcal{L}_{\text{stealthy}}\leftarrow$ distance between backdoor image $\boldsymbol{x}^{\prime}$ and clean image $\boldsymbol{x}$, $\forall\boldsymbol{x}^{\prime}\in{\mathcal{D}_{s}}^{{}^{\prime}},\forall\boldsymbol{x}\in\mathcal{D}_{s}$
7:         $\triangleright$ Equation 4
8:         $\mathcal{L}_{\text{ours}}=\mathcal{L}_{\text{stealthy}}+\mu\cdot\mathcal{L}_{\text{disentangle}}$
9:         $\phi_{\mathcal{I}}=\phi_{\mathcal{I}}-lr\cdot\frac{\partial\mathcal{L}_{\text{ours}}}{\partial\phi_{\mathcal{I}}}$
@@ D.2 Pre-training Backdoor Injector
Table 11: The detailed network structure of backdoor injector.
D.3 Backdoor Injection
The subsequent phase introduces a backdoor into the pre-trained encoder. This is achieved through the use of Shadow Dataset, which constitutes a smaller subset of the pre-training dataset. Consistent with preceding studies [29], this shadow dataset comprises 50,000 randomly selected training samples. Additionally, we employ the same reference samples as those used in BadEncoder [29]. The backdoor is injected using a designated algorithm, which leverages a combination of $\mathcal{L}_{consistency}$, $\mathcal{L}_{utility}$, and $\mathcal{L}_{alignment}$. The notations are the same as the main paper.
@@ D.3 Backdoor Injection
$\mathcal{L}_{consistency}=-\frac{\sum_{i=1}^{t}\sum_{j=1}^{r_{i}}s\left(\mathcal{F}_{\theta}^{\prime}\left(\boldsymbol{x}_{ij}\right),\mathcal{F}_{\theta}\left(\boldsymbol{x}_{ij}\right)\right)}{\sum_{i=1}^{t}r_{i}}$ (6)
@@
$\mathcal{L}_{utility}=-\frac{1}{\left|\mathcal{D}_{s}\right|}\cdot\sum_{\boldsymbol{x}\in\mathcal{D}_{s}}s\left(\mathcal{F}_{\theta}^{\prime}(\boldsymbol{x}),\mathcal{F}_{\theta}(\boldsymbol{x})\right)$ (7)
D.4 Training Downstream Classifiers
@@
D.5 Technical Background
HSV Color Space. The HSV color model describes color based on three primary attributes: hue, saturation, and value. Hue (H) represents the color itself, essentially the aspect of color we typically refer to by names like red or yellow. Saturation (S) indicates the vividness or richness of the color. A higher saturation means a more intense, pure color. Value (V), often referred to as brightness, measures the lightness or darkness of the color [67]. HSV color space can be converted from RGB color space through Equation 8, Equation 9, Equation 10, Equation 11 [63].
HSL Color Space. The HSL color model is defined by three cylindrical coordinates: hue (H), saturation (S), and lightness (L), which together capture the subtleties of color. HSL color space can be converted from RGB color space through Equation 8, Equation 9, Equation 12, Equation 13 [63].
$\mathrm{R}^{\prime}=\mathrm{R}/255;$ (8)
@@ D.5 Technical Background
$\mathrm{B}^{\prime}=\mathrm{B}/255$
$\mathrm{C}_{\max}=\operatorname{MAX}\left(\mathrm{R}^{\prime},\mathrm{G}^{\prime},\mathrm{B}^{\prime}\right)$
$\mathrm{C}_{\min}=\operatorname{MIN}\left(\mathrm{R}^{\prime},\mathrm{G}^{\prime},\mathrm{B}^{\prime}\right)$
$\Delta=\mathrm{C}_{\max}-\mathrm{C}_{\min}$
$\mathrm{H}=\left\{\begin{array}[]{l}60^{\circ}\times\left(\frac{\mathrm{G}^{\prime}-\mathrm{B}^{\prime}}{\Delta}\bmod 6\right),\mathrm{C}_{\max}=\mathrm{R}^{\prime}\\ 60^{\circ}\times\left(\frac{\mathrm{B}^{\prime}-\mathrm{R}^{\prime}}{\Delta}+2\right),\mathrm{C}_{\max}=\mathrm{G}^{\prime}\\ 60^{\circ}\times\left(\frac{\mathrm{R}^{\prime}-\mathrm{G}^{\prime}}{\Delta}+4\right),\mathrm{C}_{\max}=\mathrm{B}^{\prime}\end{array}\right.$ (9)
$\mathrm{S}=\left\{\begin{array}[]{cc}0\quad,&\Delta=0\\ \frac{\Delta}{\mathrm{C}_{\max}},&\Delta\neq 0\end{array}\right.$ (10)
$\mathrm{V}=\mathrm{C}_{\max}$ (11)
$\mathrm{S}=\left\{\begin{array}[]{cl}0,&\Delta=0\\ \frac{\Delta}{1-|2\mathrm{L}-1|}&,\Delta\neq 0\end{array}\right.$ (12)
$\mathrm{L}=\left(\mathrm{C}_{\max}+\mathrm{C}_{\min}\right)/2$ (13)
@@
D.6 Experiments Details
D.6.1 Dataset Details
We use the following datasets in our method evaluation.
@@ D.6.1 Dataset Details
D.6.2 Evaluation Metrics Details
The meanings of evaluation metric are shown below:
  • CA: This is the accuracy of a clean downstream classifier in correctly classifying the clean testing images from the corresponding downstream dataset.
  • BA: BA measures the accuracy of a backdoored downstream classifier on the same clean testing images from the corresponding downstream dataset.
  • ASR: ASR measures the percentage of these backdoored test images that are classified as the target class by the backdoored downstream classifier.
D.6.3 Parameter Setting Details
We adopt the default parameter settings as follows unless stated otherwise:
@@ D.6.3 Parameter Setting Details
    We set $\lambda_{1}=10,\lambda_{2}=0.025$ in the loss equation term Equation 4, $\alpha=0.1,\beta=5$ in Algorithm 1, and $\mu=0.1$ in Algorithm 2 to scale the loss terms to the same 0-1 range.
@@
    The Shadow dataset consists of 50000 images randomly chosen from the pre-training dataset.
@@
    Since we know the SSL’s pre-training process, the SSL augmentation transforms are also known. By default, the augmentation transforms used are RandomResizeCrop, RandomHorizontalFlip, ColorJitter, and RandomGrayscale.
  5. 5.
  6. 6.
    Algorithm 1 and Algorithm 2 utilize the Adam optimizer with 200 training epochs, a batch size of 256, and learning rates of 0.001 and 0.005 respectively.
D.6.4 Experimental Setup in Real World Case ImageNet
We selected a random 1% subset of ImageNet’s training images to serve as the shadow dataset. Additionally, we adjusted the size of each image in both the shadow and downstream datasets to dimensions of 224×224×3 as Google does in the encoder pre-training stage [5]. We set $\alpha=0.1,\beta=20$ in Algorithm 1, and $\mu=0.1$ in Algorithm 2 to scale the loss terms to the same range. Moreover, Algorithm 1 and Algorithm 2 here use a batch size of 16, and learning rates of $10^{-4}$ and 0.005 respectively. Other experiment settings are the same as those in Appendix D.6.3.
Appendix E Ethical Statements
This research is focused on identifying potential weaknesses in SSL that could be exploited by backdoor attacks. Our approach is guided by ethical standards, and we intend to enhance, not compromise, system security. We aim to promote awareness and encourage the creation of strong protective measures by demonstrating current susceptibilities within SSL frameworks.
Appendix F Limitation
The limitation of the current research is its exclusive focus on computer vision (CV) models. This specificity might not address potential vulnerabilities or behaviors applicable to other types of models, such as those used in natural language processing (NLP) or audio signal processing.
Appendix G Broader Impacts
For potential positive societal impacts, first by exposing potential vulnerabilities in SSL models, the research encourages the development of more robust security measures. This can lead to safer deployment of AI systems in critical applications such as healthcare, autonomous vehicles, and financial services, where security is paramount. Moreover, Highlighting the vulnerabilities in AI systems can spur more research into ethical AI practices. This includes developing guidelines and frameworks to ensure that AI technologies are used responsibly and do not inadvertently harm users or perpetuate biases.
For potential negative societal impacts, public knowledge of vulnerabilities in AI systems can lead to a loss of trust in these technologies. If people believe that AI systems can be easily manipulated or compromised, they may be hesitant to rely on these technologies, slowing down technological adoption and potential benefits.

### v2

@@ 6 Conclusions and Future Work
Grad-CAM. We assess INACTIVE’s resilience to Grad-CAM [58], which generates a heatmap for a given model and input sample, where the heat intensity of each pixel reflects its significance in the model’s final prediction. Grad-CAM is effective for identifying smaller Trojans [41, 17], as these Trojans tend to produce high heat values in compact trigger zones, leading to abnormal heatmaps. Nevertheless, our backdoor transformation function alters the entire image, rendering Grad-CAM ineffective in detecting it. Figure A2 illustrates the comparison of heatmaps from both clean images and backdoored images created by our method. The similarity in these heatmaps suggests that INACTIVE can withstand defenses based on Grad-CAM.
@@
Various Noises. Next, we are going to introduce the noises, i.e., JPEG compression, Salt&Pepper noise, and Poisson noise, applied in the robustness test.
@@
Figure A3: Resilience to JPEG compression, Salt&Pepper noise, Poisson noise of different densities.
@@
Evaluating the robustness of our method across these noises helps ensure its performance is reliable even when image quality is compromised when encountering real-world image issues. We illustrate the evaluation results in Figure A3, where the pre-trained dataset is CIFAR10 and the downstream dataset is STL10. The results prove that our method, which utilizes $\mathcal{L}_{\text{disentangle}}$, maintains high ASRs in most cases, and the ASR is significantly higher than the cases ablating $\mathcal{L}_{\text{disentangle}}$ in our method. For example, under different intensities of Poisson noise, the ASR of our method maintains nearly 100%, while the ASR of the cases ablating $\mathcal{L}_{\text{disentangle}}$ in our method is unstable, ranging from 12% to 61%. The outcome indicates that the $\mathcal{L}_{\text{disentangle}}$ plays a prominent role in promoting the robustness of our method. The reason may be that the $\mathcal{L}_{\text{disentangle}}$ encourages the model to learn features that remain invariant under the transformations. Namely, the model can still recognize the embedded triggers even when various noises or other image perturbations are applied. This aspect of feature preservation is key to ensuring that the backdoor attack remains effective regardless of the image transformations or certain noises applied.
Figure A4: Resilient to STRIP [14]. INACTIVE can withstand the STRIP runtime defense.
@@
STRIP. STRIP [14] scrutinizes a suspect sample by overlaying diverse image patterns onto it and monitoring the uniformity of the model’s predictions for these altered inputs. A low entropy score, suggesting consistent predictions across these perturbed samples, would lead STRIP to classify the sample as Trojan-infected. The experimental outcomes depicted in Figure A4 reveal that the entropy values for clean and Trojan-compromised models produced through our method overlap considerably, signifying that our attack can withstand the STRIP runtime defense. INACTIVE manages to evade STRIP as the superimposition process destroys the attack’s trigger, thereby causing both the Trojan-induced and clean samples to exhibit significant alterations in prediction when superimposed, aligning them with what is observed in clean cases.
Adaptive Defense.  Our threat model assumes the attacker controls pre-downstream training, while the defender can only intervene during downstream training and inference. Since our trigger injection operates in the HSV color space, we introduce Salt and Pepper Noise and Poisson Noise perturbations in the HSV color space of both downstream training and inference phase, simulating potential adaptive defense strategies a defender might use to disrupt the attack. We use CIFAR10 as the pre-training dataset and GTSRB, STL10, and SVHN as downstream datasets. Experimental results show INACTIVE’s ASR remains unaffected, demonstrating its robustness. This resilience stems from our disentanglement loss, which increases the distributional gap between backdoor and augmented samples. This design inherently enhances the robustness of the trigger, as the larger distributional gap in the color space ensures that the trigger remains distinguishable even when noise is introduced to the samples. Consequently, the trigger is less affected by noisy samples and retains its effectiveness in activating the backdoor.
Downstream dataset Salt and Pepper on HSV Poisson Noise on HSV
BA ASR BA ASR
GTSRB 81.96% 100% 81.81% 100%
STL10 73.81% 99.38% 73.65% 99.41%
SVHN 60.17% 95.34% 60.65% 91.11%
Table A7: Adaptive defense evaluation results pre-trained on CIFAR10.
A5.2 Sensitivity Evaluation
@@
Figure A5 illustrates the sensitivity evaluation of the hyper-parameters alpha and beta, showcasing that the model’s performance exhibits relative insensitivity to parameter variations. Specifically, for alpha, both the Attack Success Rate (ASR) and Backdoored Accuracy (BA) remain relatively constant across a range of alpha values from 0.05 to 0.30, suggesting that once an optimal threshold is reached, the model’s performance is stable despite further adjustments to alpha. Similarly, for beta, changes across a broader spectrum, from 0 to 10, show that ASR quickly stabilizes at 100%, while BA and CA display minimal fluctuations, indicating the robust model performance that is not significantly affected by variations in beta. The horizontal axis in both graphs represents the incremental values of alpha and beta, demonstrating the model’s consistent performance over these parameter ranges.
Figure A5: Sensitivity evaluation of hyper-parameters. INACTIVE performs consistently over these parameter ranges.
A5.3 Results Analysis
@@
Figure A6: Visualization of feature space segregation over epochs demonstrates the efficacy of our backdoor trigger in creating a distinguishable separation between backdoor and clean features, unlike WaNet [47] where the features remain intermingled, highlighting our method’s utility in securing encoders against backdoor threats.
@@
Figure A6 presents a visual analysis of the feature shifts induced by backdoor injection across different epochs for our proposed method compared to the WaNet [47]. Throughout the backdoor injection process, it becomes evident that our method’s trigger effectively separates the backdoor features from the clean features. This distinction is observable as the training progresses from epoch 10 through to epoch 200, where the separation between the backdoor and clean features becomes increasingly pronounced, indicating a clear demarcation that our encoder can exploit to differentiate between the two.
@@
[…]
Epoch 10 20 50 100
[…]
Table A8: KL divergence values across different epochs.
A5.4 Stealthiness Evaluation
A5.4.1 Algorithmic Metrics
Table A9 and Table A10 show the detailed experimental results of the stealthy metrics across various backdoor attack methods and datasets.
[…]
Badencoder [27] /DRUPE [68] CIFAR10 STL10 0.8375 14.3745 0.03894 0.7788 63.7301
SVHN 0.8380 15.0739 0.11092 0.8518 33.1811
GTSRB 0.8309 12.8845 0.08094 0.8295 63.1784
STL10 CIFAR10 0.8455 14.9113 0.03696 0.8598 13.8215
SVHN 0.8407 15.0849 0.11110 0.8165 64.5712
GTSRB 0.8308 12.8965 0.08597 0.8306 63.2014
Average 0.8372 14.2043 0.07747 0.8278 50.2806
CTRL [34] CIFAR10 STL10 0.9415 32.4195 0.00017 0.9115 47.7871
SVHN 0.8695 32.3533 0.00065 0.7032 157.6731
GTSRB 0.8965 32.4565 0.00021 0.9787 7.9540
STL10 CIFAR10 0.9356 32.4603 0.00019 0.9827 11.9283
SVHN 0.8721 32.4198 0.00062 0.9347 59.4134
GTSRB 0.8945 32.5226 0.00024 0.9773 8.8889
Average 0.9016 32.4387 0.00035 0.9147 48.9408
WaNet [47] CIFAR10 STL10 0.7378 14.3472 0.04657 0.6611 97.1471
SVHN 0.7941 15.2503 0.10023 0.5927 129.6770
GTSRB 0.7792 13.1140 0.07615 0.7320 67.4510
STL10 CIFAR10 0.7728 14.9430 0.03996 0.6510 96.0399
SVHN 0.8026 15.4044 0.09544 0.5924 129.8306
GTSRB 0.7687 13.1645 0.08358 0.7314 67.7673
Average 0.7759 14.3706 0.07366 0.6601 97.9855
Ins-Kelvin [42] CIFAR10 STL10 0.4427 15.7024 0.11110 0.7011 107.4607
SVHN 0.4213 15.4081 0.20070 0.5905 121.0396
GTSRB 0.6224 17.4669 0.10820 0.7393 60.8467
STL10 CIFAR10 0.4624 15.6372 0.10970 0.6582 85.2866
SVHN 0.4358 15.4436 0.19840 0.6342 130.1333
GTSRB 0.5942 17.5105 0.11640 0.7422 62.0004
Average 0.4965 16.1948 0.14075 0.6776 94.4612
Ins-Xpro2 [42] CIFAR10 STL10 0.5672 17.8706 0.03451 0.7806 66.9640
SVHN 0.5324 17.4862 0.06699 0.7698 27.4974
GTSRB 0.6947 18.3951 0.03151 0.8997 10.7916
STL10 CIFAR10 0.5354 17.9291 0.03806 0.8063 19.0678
SVHN 0.5057 17.4956 0.07152 0.7505 45.3211
GTSRB 0.6541 18.4070 0.03632 0.8932 11.9285
Average 0.5816 17.7640 0.04615 0.8167 30.2617
POIENC [39] CIFAR10 CIFAR10 0.1214 11.2787 0.15867 0.5967 172.2200
BLTO [63] CIFAR10 CIFAR10 0.8417 29.6756 0.00941 0.9501 36.3848
SSLBKD [56] CIFAR10 CIFAR10 0.8737 16.2414 0.09640 0.8913 118.3200
Ours CIFAR10 STL10 0.9405 24.0804 0.02311 0.9161 45.8759
SVHN 0.9807 46.0749 0.00040 0.9976 1.8810
GTSRB 0.9687 37.4394 0.00338 0.9944 1.2025
STL10 CIFAR10 0.9928 46.5797 0.00019 0.9981 0.5018
SVHN 0.9847 45.7410 0.00031 0.9464 30.0092
GTSRB 0.9905 46.5149 0.00011 0.9977 0.2157
Average 0.9763 41.0717 0.0046 0.9751 13.2810
Table A9: Stealthiness Metrics across Methods, Pre-training and Downstream Datasets.
Model Pre-training Dataset Downstream Dataset SSIM PSNR(dB) LPIPS
ISSBA [36] ImageNet STL10 0.7836 30.3333 0.05615
GTSRB 0.7292 31.7344 0.11651
SVHN 0.6860 31.9810 0.20006
Average 0.7329 31.3496 0.12424
Ours ImageNet STL10 0.9900 38.3300 0.00400
GTSRB 0.9900 32.1400 0.01700
SVHN 0.9800 33.2500 0.01600
Average 0.9867 34.5733 0.01233
Table A10: Stealthiness Metrics across Methods on ImageNet.
A5.4.2 Human Inspection
We conducted a human inspection study [47] to evaluate the stealthiness of various backdoor attacks. We introduce the detailed settings in section A6. The results in Table A11 indicate that traditional methods like SSLBKD, BadEncoder/DRUPE, and POIENC have relatively low fooling rates, with SSLBKD at just 4.9% overall, suggesting they are easier for humans to detect. In contrast, our method achieves a higher fooling rate of 49.5% overall, indicating much greater stealth. However, as shown in Figure A9, subtle artifacts, such as slight color changes, can still be detected by humans.
Images SSLBKD CTRL BadEncoder /DRUPE POIENC Ins-Kelvin BLTO ISSBA Ours
Backdoor 2.4 31.2 4.0 9.6 12.8 22.4 24.8 50.6
Clean 7.4 27.2 8.8 16.8 10.4 27.6 26.4 48.4
Both 4.9 29.2 6.4 13.2 11.6 25.0 25.6 49.5
Table A11: Success fooling $\uparrow$ (%) of different methods. Our method achieves the greatest stealth with the highest overall success fooling rates.
A6 Implementation Details
The following sections provide a comprehensive breakdown of the process of implementing the attack, which is methodically organized into three distinct phases: pre-training the encoder, injecting the backdoor, and training the downstream classifiers. Additionally, we explain the HSV&HSL Color Spaces and Wasserstein Distance used in the method in detail. Moreover, we provide more details on the datasets and experimental setup details.
A6.1 Pre-training Image Encoders
@@
Table A12: The detailed network structure of backdoor injector.
A6.2 Backdoor Injection
The subsequent phase introduces a backdoor into the pre-trained encoder. This is achieved through the use of Shadow Dataset, which constitutes a smaller subset of the pre-training dataset. Consistent with preceding studies [27], this shadow dataset comprises 50,000 randomly selected training samples. Additionally, we employ the same reference samples as those used in BadEncoder [27]. The backdoor is injected using a designated algorithm, which leverages a combination of $\mathcal{L}_{\text{consistency}}$, $\mathcal{L}_{\text{utility}}$, and $\mathcal{L}_{\text{alignment}}$. The notations are the same as the main paper.
@@
$\mathcal{L}_{\text{consistency}}=-\frac{\sum_{i=1}^{t}\sum_{j=1}^{r_{i}}s\left(\mathcal{F}_{\theta}^{\prime}\left(\boldsymbol{x}_{ij}\right),\mathcal{F}_{\theta}\left(\boldsymbol{x}_{ij}\right)\right)}{\sum_{i=1}^{t}r_{i}}$ (A1)
@@
$\mathcal{L}_{\text{utility}}=-\frac{1}{\left|\mathcal{D}_{s}\right|}\cdot\sum_{\boldsymbol{x}\in\mathcal{D}_{s}}s\left(\mathcal{F}_{\theta}^{\prime}(\boldsymbol{x}),\mathcal{F}_{\theta}(\boldsymbol{x})\right)$ (A2)
A6.3 Training Downstream Classifiers
@@
A6.4 Technical Background
HSV Color Space. HSV and HSL color spaces align more closely with human color perception, unlike RGB’s additive mixing. They allow for intuitive adjustments of tint, shade, and tone through single-dimension modifications. They have wide applications in the computer vision domain [38, 12], as detailed in subsection A6.4.
The HSV color model describes color based on three primary attributes: hue, saturation, and value. Hue (H) represents the color itself, essentially the aspect of color we typically refer to by names like red or yellow. Saturation (S) indicates the vividness or richness of the color. A higher saturation means a more intense, pure color. Value (V), often referred to as brightness, measures the lightness or darkness of the color [61]. HSV color space can be converted from RGB color space through Equation A3, Equation A4, Equation A5, Equation A6 [57].
HSL Color Space. The HSL color model is defined by three cylindrical coordinates: hue (H), saturation (S), and lightness (L), which together capture the subtleties of color. HSL color space can be converted from RGB color space through Equation A3, Equation A4, Equation A7, Equation A8 [57].
$\mathrm{R}^{\prime}=\mathrm{R}/255;$ (A3)
@@
$\mathrm{B}^{\prime}=\mathrm{B}/255;$
$\mathrm{C}_{\max}=\operatorname{MAX}\left(\mathrm{R}^{\prime},\mathrm{G}^{\prime},\mathrm{B}^{\prime}\right);$
$\mathrm{C}_{\min}=\operatorname{MIN}\left(\mathrm{R}^{\prime},\mathrm{G}^{\prime},\mathrm{B}^{\prime}\right);$
$\Delta=\mathrm{C}_{\max}-\mathrm{C}_{\min};$
$\mathrm{H}=\left\{\begin{array}[]{l}60^{\circ}\times\left(\frac{\mathrm{G}^{\prime}-\mathrm{B}^{\prime}}{\Delta}\bmod 6\right),\quad\mathrm{C}_{\max}=\mathrm{R}^{\prime}\\ 60^{\circ}\times\left(\frac{\mathrm{B}^{\prime}-\mathrm{R}^{\prime}}{\Delta}+2\right),\quad\mathrm{C}_{\max}=\mathrm{G}^{\prime}\\ 60^{\circ}\times\left(\frac{\mathrm{R}^{\prime}-\mathrm{G}^{\prime}}{\Delta}+4\right),\quad\mathrm{C}_{\max}=\mathrm{B}^{\prime}\end{array}\right.$ (A4)
$\mathrm{S}=\left\{\begin{array}[]{cc}0\quad,&\Delta=0\\ \frac{\Delta}{\mathrm{C}_{\max}},&\Delta\neq 0\end{array}\right.$ (A5)
$\mathrm{V}=\mathrm{C}_{\max}$ (A6)
$\mathrm{S}=\left\{\begin{array}[]{cl}0,&\Delta=0\\ \frac{\Delta}{1-|2\mathrm{L}-1|}&,\Delta\neq 0\end{array}\right.$ (A7)
$\mathrm{L}=\left(\mathrm{C}_{\max}+\mathrm{C}_{\min}\right)/2$ (A8)
@@
A6.5 Experiments Details
Referece images. Following BadEncoder [27], the reference images used in the experiments are shown in Figure A7 and Figure A8.
(a) CIFAR10
(b) SVHN
(c) GTSRB
(d) STL10
Figure A7: The default reference inputs for CIFAR10 SVHN, GTSRB, and STL10.
(a) SVHN
(b) GTSRB
(c) STL10
Figure A8: The reference inputs for attacking the image encoder pre-trained on ImageNet by Google.
Dataset Details.  We use the following datasets in our method evaluation.
@@
Parameter Setting Details.  We consider a scenario where the attacker chooses a specific downstream task/dataset, a specific target class, and a specific reference input. For the datasets CIFAR10, STL10, GTSRB, and SVHN, the selected target classes are “airplane”, “truck”, “priority sign”, and “digit one”, respectively. CIFAR10 and STL10 are used as the default pre-training datasets. The shadow dataset following [27] consists of 50000 images randomly chosen from the pre-training datasets. We adopt the default parameter settings as follows unless stated otherwise:
@@
    We set $\lambda_{1}=10,\lambda_{2}=0.025$ in the loss equation term Equation 5, $\alpha=0.1,\beta=5$ in Alg. 1, and $\mu=0.1$ in Alg. 2 to scale the loss terms to the same 0-1 range.
@@
    Alg. 1 and Alg. 2 utilize the Adam optimizer with 200 training epochs, a batch size of 256, and learning rates of 0.001 and 0.005 respectively.
Experimental Setup in Multi-modal Model. Due to the unavailability of CLIP’s original pre-training dataset, we utilize the training images from CIFAR-10, resized to 224×224×3, as our shadow dataset. We define $\alpha=0.1$ and $\beta=20$ in Alg. 1, and $\mu=0.1$ in Alg. 2 to normalize the loss terms across similar scales. Additionally, Alg. 1 and Alg. 2 are configured with a batch size of 16, and learning rates are set at $10^{-6}$ and 0.005, respectively, for a duration of 100 epochs. SVHN serves as the downstream dataset with the target label “one”. We employ the same reference inputs as described in [27]. All other experimental settings remain consistent with those outlined in the Parameter Setting section of the main paper.
Experimental Setup in Real World Case ImageNet.  We selected a random 1% subset of ImageNet’s training images to serve as the shadow dataset. Additionally, we adjusted the size of each image in both the shadow and downstream datasets to dimensions of 224×224×3 as Google does in the encoder pre-training stage [6]. We set $\alpha=0.1,\beta=20$ in Alg. 1, and $\mu=0.1$ in Alg. 2 to scale the loss terms to the same range. Moreover, Alg. 1 and Alg. 2 here use a batch size of 16, and learning rates of $10^{-4}$ and 0.005 respectively.
Experimental Setup in Human Inspection Study. For each question, we randomly selected 50 images from the ImageNet dataset, generated backdoor variants for each attack method, and listed them in random order. Each method’s set included both original and backdoor images, totaling 100 images. Before the survey, we thoroughly briefed each participant on backdoor triggers and baseline methods, confirming their understanding. We also manually reviewed each survey response for completeness.
To reduce bias, we included participants of diverse genders and ages. In all questions, images were presented in random order. Five trained participants then assessed each image, identifying it as either a backdoor or a clean sample.
To ensure consistency, we manually reviewed any outlier responses and asked participants additional questions on backdoor basics to confirm their task understanding.
Figure A9: More example ImageNet images with triggers generated by INACTIVE. The backdoor image is on the left side of the same set of images and the clean image is on the right.